Morgan Paisley is single and works part-time in a Nashville, Tennessee Hotel as a nonexempt housekeeper. Morgan gets paid $14.50 per hour and is paid on a biweekly basis. During the pay period, she worked 52 regular hours, with 12 of those considered overtime hours. She has one withholding and no pretax deductions. Morgan will pay $58.00 in federal taxes.

Required: What is her net pay? Round calculations to two decimal places and share calculations.

Employee Names Married/Single # of Withhold Regular Wages OT Wages Gross Federal WH Social Security WH Medicare WH Net Pay
Answer #1
Employee Names Married/Single # of Withhold Regular Wages OT Wages Gross Federal WH Social Security WH 6.20% Medicare WH 1.45% Net Pay
Morgan Paisley Single 1 $580.00 $261.00 $ 841.00 $ 58.00 $ 52.14 $ 12.19 $ 718.67
Employee Names Married/Single # of Withhold Regular Wages OT Wages Gross Federal WH Social Security WH 6.20% Medicare WH 1.45% Net Pay
Morgan Paisley Single 1 =40*14.5 =12*14.5*1.5 =580+261 58 =841*6.2% =841*1.45% =841-58-52.14-12.19